OpenStudy (anonymous):

Use equations to solve the following questions Daniel and Travis are twins, but their sister Phillipa is 3 years older. If the sum of their three ages is 36, how old are the twins?

OpenStudy (anonymous):

You know how old Daniel and Travis are? ^.^

OpenStudy (anonymous):

Of course we don't that's why it's being asked :D So let's say they're both x years old, okay? Just tell me when you're ready to jump in ^.^

Directrix (directrix):

Let D = T (twins have same age) Also, T = D Let P = D + 3 D + T + P = 36 D + D + D + 3 = 36 Solve for D @mysteryman123

OpenStudy (anonymous):

Who cares about age anyway? ^.^

OpenStudy (anonymous):

Let D and T be x, so P=(x+3) Therefore,x+x+(x+3)=36 3x=33 x=11 The twins are 11 year old